OpenStudy (anonymous):

A 62 g sample of carbon has a density of 2.3 g/cm3. What is the volume of the sample?

OpenStudy (rushwr):

u got it right?

OpenStudy (anonymous):

For the first question I got the answer, but i'm still confused.

OpenStudy (rushwr):

Use \[Density = \frac{ mass }{ Volume }\] Here we know the mass and the density So: \[Volume = \frac{ mass }{ Density }\]

OpenStudy (rushwr):

\[Volume = \frac{ 62g }{ 2.3gcm ^{-3} }\]

OpenStudy (rushwr):

Got it ? It's just the matter of subjecting the unknown !

OpenStudy (anonymous):

thanks so much :) @Rushwr

OpenStudy (rushwr):

Now do u understand ?

OpenStudy (anonymous):

Yes.

OpenStudy (rushwr):

Gud ~~~~ :)